A ripple tank produces water waves through the up and down movement of a wooden bar. The bar moves up 20 times in 4 s. The distance between each water wave is 0.04 m 


  1. Calculate  the frequency of the waves.
  2. State the wavelength of the waves.

            

3.Calculate the speed of the waves


4.The ripple tank is 0.8m long. Calculate the time it takes a water wave to travel from one end to the other. 



1
Expert's answer
2021-02-28T07:40:05-0500

1)


"f=\frac{20}{4}=5\ Hz"

2)


"\lambda=0.4\ m"

3)


"v=\lambda f=(0.4)(5)=2\frac{m}{s}"

4)


"t=\frac{l}{v}=\frac{0.8}{2}=0.4\ s"